OpenStudy (anonymous):

I don't know why I keep getting this question wrong. Can someone help please? Identify the 16th term of a geometric sequence where a1 = 4 and a8 = -8,748 (A) -172,186,884 (B) -57,395,628 (C) 57,395,628 (D) 172,186,884

OpenStudy (anonymous):

\[a_n = a_1*r ^{n-1}\]

OpenStudy (anonymous):

Find the ratio between each term, then use that equation, using n=16

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} term&value \\\hline\\ a_1&4\\ a_2&4r\\ a_3&(4r)r\\ a_4&(4rr)r\\ a_5&(4rrr)r\\ a_6&(4rrrr)r\\ a_7&(4rrrrr)r\\ a_8&(4rrrrrr)r=-8748 \end{array} \) what does that give you for the commond ratio "r"?

OpenStudy (anonymous):

I'm sorry, i do not understand.

OpenStudy (anonymous):

I am not used to using "r"

OpenStudy (jdoe0001):

hmm do you know what a geometric sequence is?

OpenStudy (anonymous):

Yes, I do.

OpenStudy (jdoe0001):

well... common ratio "r" is just the multiplier we dunno what the multiplier is, to get the next term but we do know that, it's used each time, every time to get the "next term" notice above so.... notice the last term there, the 8th one "r" was used 7 times to get there, that is, to get -8748 so.... solving for "r", that gives us?

OpenStudy (jdoe0001):

hmm got a minus there for some reason... lemme fix that

OpenStudy (jdoe0001):

once you get the "r" the rest is peanuts as larryboxaplenty already pointed out \(\bf a_{\color{brown}{ n}}=a_1{\color{blue}{ r}}^{{\color{brown}{ n}}-1}\qquad {\color{brown}{ n}}=16\qquad a_{\color{brown}{ 16}}=a_1{\color{blue}{ r}}^{{\color{brown}{ 16}}-1} \)

OpenStudy (anonymous):

I think he/she is confused on finding "r"

OpenStudy (jdoe0001):

hmm pretty straightforward it seems though

OpenStudy (anonymous):

Yes. I am confused on finding "r" I am sorry.

OpenStudy (jdoe0001):

hmm ahemm one sec

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} term&value \\\hline\\ a_1&4\to& 4\\ a_2&4r\to &4r\\ a_3&(4r)r\to &4r^2\\ a_4&(4rr)r\to &4r^3\\ a_5&(4rrr)r\to &4r^4\\ a_6&(4rrrr)r\to &4r^5\\ a_7&(4rrrrr)r\to &4r^6\\ a_8&(4rrrrrr)r\to &4r^7=-8748 \end{array}\) ok... see the equation? what do you think is "r"?

OpenStudy (anonymous):

Lol, I got the answer! Thank you! Sorry, this stuff was just going right over my head. Thank you all for your help!

OpenStudy (wolf1728):

Gee, I was all set to tell you: r^7 = -8,748/4 r^7 = -2,187 r = 7th root of -2,187

OpenStudy (wolf1728):

anyway, u r welcome